# Transfer Learning with Foundational Models for Time Series Forecasting using Low-Rank Adaptations

### v1

@@ 3 LLIAM: Llama Lora-Integrated Autorregresive Model
LLMs are pre-trained with massive amounts of data from different sources. Due to the vast volume of information that a model has to process, significant computational power and time is demanded to build such an extensive knowledge. PEFT techniques enable us to rapidly adapt these models to specific tasks, leveraging their prior knowledge. Our contribution adapts and evaluates the LLM LlaMA for TSF using the LoRA reparametrization technique based on its open-source Lit-LLaMA implementation [^1] https://github.com/Lightning-AI/lit-llama. The resulting combination is called LLIAM: The Llama Lora-Integrated Autorregresive Model.
@@
As mentioned in Section 2.5, the LLaMA family is a collection of several foundational text models. Multiple versions of this model with different numbers of parameters (weights) are released, ranging between 7 billion (7B) and 65 billion (65B). Our proposal employs the 7B model because it is the most lightweight one. A stacked Transformer decoder-only architecture is employed. Before entering the information into the decoder stack, the input series is transformed into a textual prompt and converted into a numerical vector using embedding. RMS Normalization is applied before processing the data in the Multi-Headed Attention module (MHA) and the Feed Forward Network (FNN) to stabilize the training of the model. MHA allows learning the relationships between the elements of the input data. An efficient implementation of it is used, with a cache that allows the recycling of previous computations to reduce the computational cost. Finally, an FFN is used to improve the features computed by MHA. The output of one decoder is used as the input of the next. Finally, the last output projected with a linear layer, used to predict the next token of the series, which will be added to the input to infer the next one.
@@
In Figure 2 the proposed architecture of LLIAM is presented. The following paragraphs describe its main aspects.
@@
LLaMA can only process textual information as its input as a LLM. Time-series are prompted into the model, following a similar prompting scheme to Prompcast [46]. The numerical sequential data is transformed into textual strings following the pattern presented in Table 1. As seen, our proposal aims for a more generic pattern that does not introduce any exogenous variable or data into the prompt. In contrast with the approach proposed by Promptcast, only the number (n) and value of the input observations (series) and the desired forecast horizon (h) is provided to the model. The model is expected to only return the desired future values of the series ($t_{n+1},...,\hat{t}_{n+h}$) This will enable the use of the same template for all the datasets and evaluate the model output, relying on only its capabilities to extract the different patterns presented in the input series without providing any additional helpful information.
@@
An embedding is used to transform textual tokens into vectors of continuous values in order to ensure a correct computation of the input data by the model. In LLaMA-1, each vector has a dimensionality equal to 4096 [6]. This hyperparameter is known as the dimensionality of the model ($d_{m}$), and must be consistent across all the modules of LLaMA to ensure uniformity between their inputs and outputs. Therefore, all of them will also have $d_{m}$ dimensions.
@@
Layer Normalization (LN) [47] is employed to improve and stabilize the training of the baseline Transformer after each of its modules. However, LLaMA-1 developers opted for changing what normalization technique is applied and when. Specifically, they applied Root Mean Squared Normalization (RMS Norm) [48] before processing the inputs on every module. This modification was inspired by GPT3.
@@
The baseline Scaled Dot-Product Attention [17] is used in LLaMA. This mechanism is capable of learning relationships between two inputs computationally-efficiently. This procedure can be interpreted as a similarity measure between the different pair of elements that compose the input. In equation 1 is shown how this attention matrix is computed. For a given embedding of the input data $X$, three new representations are calculated, called queries ($Q$), keys ($K$) and values ($V$). Queries represent what an element of $X$ wants to attend, keys show what valuable information a token can provide and $V$ what an element really is. Each of these representations are generated by multiplying three learnable weights matrices to $X$. The similarity between the queries and keys is determined by their dot product, represented as $QK^{T}$. Higher values indicate greater similarity. A softmax function is applied to produce a probability distribution that describes how much attention each query pays to each key.
@@
The autoregressive nature of LLaMA leads to repeating redundant calculations that remain unchanged when adding the predicted output tokens to the input. To speed up the computation of MHA, the queries and keys of the input sentence are cached at each step. When predicting the $i^{th}$ token, two auxiliary matrices are used to retrieve these representations from the first to the $(i-1)^{th}$ token. Consequently, only the key and value for the $i^{th}$ element need to be computed instead of the full $Q$ and $V$ matrices. Also, these new values are stored in these caches, facilitating the calculation for the $(i+1)^{th}$ token.
@@
Instead of using the absolute positional encoding proposed by Vaswani et al. [17], LLaMA employs a relative positional encoding method known as RoPE [49]. This technique integrates positional information into the MHA module using a 2D rotation matrix, applicable to any even-dimensional space. In essence, for an element of the input at position $m$, RoPE rotates the corresponding key and query by $m$ times a fixed angle $\theta$. When the dimensionality of the keys ($d_{K}$) and queries ($d_{V}$) are higher than two, dimensions are grouped in $d/2$ pairs and rotated following the set of different rotation angles ($\Theta$) given by equation 3.
@@
FFNs are used at the end of each decoder to introduce non-linearity while the MHA module processes relationships between positions. The original Transformer architecture uses an FFN with a ReLU activation function. However, LLaMA-1 improves performance by using the SwiGLU activation function [50], which has been shown to outperform Transformers using ReLU on various language understanding tasks.
@@
As introduced in Section 2.4, Low-Rank Adaptation (LoRA) [7] is a PEFT technique used for fine-tuning FMs such as LLaMA. This method can be incorporated into the model during production deployment, resulting in no additional inference latency, unlike other techniques such as adapters. LoRA was motivated by the conclusions obtained in [51], which specify that as models become larger, fewer parameters are needed to fine-tune them to achieve a certain accuracy on a specific task (a concept known as intrinsic dimension). Therefore, the change in the parameters of a model when fine-tuning may also have a low intrinsic dimension. A reparameterization such as LoRA may achieve competitive results in LLMs.
Equation LABEL:eq:lora-base describes how the LoRA technique is added into a DL module. Given a pretrained model with $W^{model}\in\mathcal{R}^{d_{m}\times k}$ weights, the change needed for fine-tuning it to a certain task is given by $\Delta W$. Based on the previous hypothesis, it must have a low intrinsic dimension. This is ensured by expressing $\Delta W$ as a low-rank decomposition of two matrices $A$ and $B$, where $A\in\mathcal{R}^{r\times k}$ and $B\in\mathcal{R}^{d_{model}\times r}$. $r$ is the hyperparameter responsible for defining the maximum desired rank of these matrices, therefore it must be $r\ll min(k,d)$. It should be noted that $W^{model}$ is not changing during the fine-tuning phase, and only $A$ and $B$ are trainable. The matrix $B$ is initialized to 0 to ensure that $\Delta W=0$ at the beginning of the training.
@@
LoRA can be applied to almost every module of LLaMA, but we have only employed it over the queries and values projections, as illustrated in figure 2.
@@
Since LLIAM is based on the LLaMA-1 7B model, most of its hyperparameters are identical to those detailed in [6]. Table 2 lists the main hyperparameters of LLIAM. #heads, #embd, and #decoders refer to the base model architecture, indicating the number of heads in the MHA, the embedding dimension, and the number of stacked decoders, respectively. The remaining hyperparameters relate to the LoRA architecture and training. With $r$ set to 8 and $\alpha$ set to 16, the scale factor when merging the model and LoRA outputs is 2. A light dropout is applied to the LoRA weights to prevent overfitting. The optimizer algorithm used is AdamW [52]. A micro-batch training strategy is adopted and a maximum number of training iterations is also defined. To train the LoRAs a supervised procedure is followed. The generated prompt is tokenized and transformed into a nominal representation using the SentencePieces method [53]. Then, input and target tokens are separated. The target tokens logits and the model generated logits are used to calculate a cross-entropy loss.
@@ 4 Experimentation framework
The main purpose of our experimentation is to demonstrate the capabilities of our LLM-based proposal, differencing two main groups of experiments. In the first group, the performance of LLIAM is compared between specific non-foundational DL methods (such as RNNs and TCNs, described in Section 2.1). Relative and absolute quality measures are computed to show the error committed by each model and their training times. This group of experiments is called “Experimental Study”.
The second group of experiments aims to show that our proposal also has a high generalization capacity. This is because it is able to tackle problems and domains never seen before, thanks to the exploitation of TL, a fundamental feature of FMs and therefore of LLMs. A series of tests have been proposed with datasets that have never been used before. The models have not been fitted with them either, so they face unknown domains and distributions of the data. This makes it possible to check whether the knowledge gain of the adapted model is significant with respect to the general knowledge of the base model, expecting in an increase of the ability to associate known concepts (in this work, series) with unknown ones. This paradigm is called Zero-Shot Learning (ZSL), so this group of experiments is called ”Zero-shot”.
This section outlines the comprehensive framework used to evaluate the proposed method. Section 4.1 details the datasets utilized and the testing procedures for LLIAM. Results from the experimental studies are presented in Section 4.2 and Zero-shot studies results are presented in Section 4.3.
4.1 Experimental seutp
The Monash University time series forecasting repository has been chosen for selecting almost every dataset used in this study. Its mission is to be the “first comprehensive repository of time series forecasts containing related time series datasets to facilitate the evaluation of global forecasting models” [54] thanks to its 30 heterogeneous sets, with different properties and coming from diverse domains. Of all the available sets, we filtered out all those where the prediction horizon was not clearly specified in the repository and were not univariate. Subsequently, we were left with only 8 main sets from this repository (Electricity, M1 Monthly, M1 Quarterly, M3 Monthly, M3 Quarterly, NN5 Daily, NN5 Weekly, San Francisco Traffic). The two Electricity Transformer Temperature (ETT) datasets [55], typically used in TSF, are also used in our study, having a total of 10 datasets employed. These datasets are described below:
@@ 4.1 Experimental seutp
    Electricity (1 Dataset): Contains the electricity consumption of 370 customers in kilowatts from 2011 to 2014. All series have the same length. Only the version whose sampling frequency is weekly has been selected.
@@
    M1 (2 Datasets): This dataset is composed of a selection of series from one of the first competitions dedicated to its prediction, called M competitions (M, because its creator is called Makridakis). It consists of 617 monthly time series, related to economics, industry, and demography. Not all series have the same number of measurements. From it, we have selected the versions with monthly and quarterly sampling frequency.
@@
    M3 (2 Datasets): Like the M1 dataset, it contains 1428 series from the third edition of the M competitions. They are again not of the same length and add finance-related phenomena to their domains. From it, we have selected the versions that present a monthly and quarterly sampling frequency.
@@
    NN5 (2 Datasets): Dataset from another competition, with 111 series collecting transactions from ATMs located in Great Britain. In this case, the series have the same length. From it, we have selected the versions with a daily and weekly sampling frequency.
@@
    San Francisco Traffic (1 Dataset): It contains 862 series collecting occupancy rates of freeways located in the San Francisco Bay Area between the years 2015 and 2016. Only the version whose sampling frequency is weekly has been selected.
@@
electricity 321 weekly 8 65 Experimental
m1 617 monthly 18 15 Experimental
m1 203 quarterly 8 5 Experimental
m3 1428 monthly 18 15 Experimental
m3 756 quarterly 8 5 Experimental
nn5 111 daily 56 9 Experimental
nn5 111 weekly 8 65 Experimental
@@
A sliding window approach is used to adapt the TSF problem as a supervised learning problem. An instance consists of two windows of observations, ordered in time and without gaps. The first represents the input values to be taken by the model ($X$) and the second the expected output ($Y$), i.e., the labels. Thanks to this restructuring, we were able to transform a series into a set of $(x,y)$ pairs compatible with LLIAM. With this conversion, a textual prompt is constructed following the template described in Section 3, leaving the datasets ready to be used by the model. Table 3 shows the input window size and the prediction horizon used to construct the sliding windows for each data set, other relevant properties, and in which study each dataset is used.
@@
RNN-LSTM 50, 100, 200 1, 2, 4 32, 64, 128 16, 32 1e-3, 1e-2 0
RNN-GRU 50, 100, 200 1, 2, 4 32, 64, 128 16, 32 1e-3, 1e-2 0
TCN 50, 100, 200 1, 2, 4 16, 32, 64 16, 32 1e-3, 1e-2 0, 0.1, 0.25
Table 4: Hyperparameters to be used in the experimentation for each non-foundational model. dim refers to the size of the hidden state in the recurrent networks and the number of channels generated by the convolutions in the TCN. bs indicates the batch size and lr the learning rate of the optimizer.
@@
The evaluation of the models is performed offline. The metrics selected are the Root Mean Square Error [56] (RMSE, equation 5), symmetric mean absolute percentage error [56] (SMAPE, equation 6).
@@
$SMAPE=\frac{2}{n}\sum_{t=0}^{n}\frac{|y_{t}-\hat{y_{t}}|}{\max(|y_{t}|+|\hat{y_{t}}|,117\cdot 10^{-8})}$ (6)

### v3

@@ 3 LLIAM: Llama Lora-Integrated Autoregressive Model
LLMs are pre-trained with massive amounts of data from different sources. Due to the vast volume of information that a model has to process, significant computational power and time is demanded to build such an extensive knowledge. PEFT techniques enable us to rapidly adapt these models to specific tasks, leveraging their prior knowledge. Our contribution adapts and evaluates the LLM LlaMA for TSF using the LoRA reparametrization technique based on its open-source Lit-LLaMA implementation [^1] https://github.com/Lightning-AI/lit-llama and a generalisation of the PromptCast [40] prompting scheme for time-series, which allows us to show how these models, using only the input data of conventional DL algorithms, perform on the TSF task. The resulting approach is called LLIAM: The Llama Lora-Integrated Autoregressive Model.
@@
Our proposal employs the LLaMA-1 7B model because it is the most lightweight one. A stacked Transformer decoder-only architecture is employed. Before entering the information into the decoder stack, the input series is transformed with our general prompting scheme into a textual prompt and converted into a numerical vector using an embedding. RMS Normalization is applied before processing the data in the Multi-Headed Attention module (MHA) and the Feed Forward Network (FNN) to stabilize the training of the model. MHA allows learning the relationships between the elements of the input data. An efficient implementation of it is used, with a cache that allows the recycling of previous computations to reduce the computational cost. Finally, an FFN is used to improve the features computed by MHA. The output of one decoder is used as the input of the next. Finally, the last output projected with a linear layer, used to predict the next token of the series, which will be added to the input to infer the next one.
@@
In Figure 2 the proposed architecture of LLIAM is presented. The following paragraphs describe its main mechanisms.
Figure 2: LLIAM architecture. Only Low-Rank adaptations applied to Query (Q) and Values (V) are trained. The rest of the models is freezed. Our prompting scheme is applied to the numerical input data to align with the textual representation used by the LLM.
@@
LLaMA-1 7B can only process textual information as its input as a LLM. Time-series are prompted into the model, following a simplification of the prompting scheme Prompcast [40]. The numerical sequential data is transformed into textual strings following the pattern presented in Table 1. As seen, our proposal aims for a more generic pattern that does not introduce any exogenous variable or data into the prompt. In contrast with the approach proposed by Promptcast, only the number (n) and value of the input observations (series) and the desired forecast horizon (h) is provided to the model. It is expected to only return the desired future values of the series ($t_{n+1},...,\hat{t}_{n+h}$). This will enable the use of the same template for all the datasets and evaluate the model output, relying on only its capabilities to extract the different patterns presented in the input series without providing any additional helpful information. Using a common prompt during the fine-tuning phase of the model helps differentiate between a TSF task and other unrelated tasks. This way, our approach can recognize when it is dealing with a forecasting task.
@@
An embedding is used to transform textual tokens into vectors of continuous values in order to ensure a correct computation of the input data by the model. In LLaMA-1 7B, each vector has a dimensionality equal to 4096 [6]. This hyperparameter is known as the dimensionality of the model ($d_{m}$), and must be consistent across all the modules of LLaMA-1 7B to ensure uniformity between their inputs and outputs. Therefore, all of them will also have $d_{m}$ dimensions.
@@
Layer Normalization (LN) [46] is employed to improve and stabilize the training of the baseline Transformer after each of its modules. However, LLaMA-1 7B developers opted for changing what normalization technique is applied and when. Specifically, they applied Root Mean Squared Normalization (RMS Norm) [47] before processing the inputs on every module. This modification was inspired by GPT3.
@@
The baseline Scaled Dot-Product Attention [15] is used in LLaMA-1 7B. This mechanism is capable of learning relationships between two inputs computationally-efficiently. This procedure can be interpreted as a similarity measure between the different pair of elements that compose the input. In equation 1 is shown how this attention matrix is computed. For a given embedding of the input data $X$, three new representations are calculated, called queries ($Q$), keys ($K$) and values ($V$). Queries represent what an element of $X$ wants to attend, keys show what valuable information a token can provide and $V$ what an element really is. Each of these representations are generated by multiplying three learnable weights matrices to $X$. The similarity between the queries and keys is determined by their dot product, represented as $QK^{T}$. Higher values indicate greater similarity. A softmax function is applied to produce a probability distribution that describes how much attention each query pays to each key.
@@
The autoregressive nature of LLaMA-1 7B leads to repeating redundant calculations that remain unchanged when adding the predicted output tokens to the input. To speed up the computation of MHA, the queries and keys of the input sentence are cached at each step. When predicting the $i^{th}$ token, two auxiliary matrices are used to retrieve these representations from the first to the $(i-1)^{th}$ token. Consequently, only the key and value for the $i^{th}$ element need to be computed instead of the full $Q$ and $V$ matrices. Also, these new values are stored in these caches, facilitating the calculation for the $(i+1)^{th}$ token.
@@
Instead of using the absolute positional encoding proposed by Vaswani et al. [15], LLaMA-1 7B employs a relative positional encoding method known as RoPE [48]. This technique integrates positional information into the MHA module using a 2D rotation matrix, applicable to any even-dimensional space. In essence, for an element of the input at position $m$, RoPE rotates the corresponding key and query by $m$ times a fixed angle $\theta$. When the dimensionality of the keys ($d_{K}$) and queries ($d_{V}$) are higher than two, dimensions are grouped in $d/2$ pairs and rotated following the set of different rotation angles ($\Theta$) given by equation 3.
@@
FFNs are used at the end of each decoder to introduce non-linearity while the MHA module processes relationships between positions. The original Transformer architecture uses an FFN with a ReLU activation function. However, LLaMA-1 7B improves performance by using the SwiGLU activation function [49], which has been shown to outperform Transformers using ReLU on various language understanding tasks.
@@
As introduced in Section 2.3, Low-Rank Adaptation (LoRA) [7] is a PEFT technique used for fine-tuning FMs. This method can be incorporated into the model during production deployment, resulting in no additional inference latency, unlike other techniques such as adapters. LoRA was motivated by the conclusions obtained in [50], which specify that as models become larger, fewer parameters are needed to fine-tune them to achieve a certain accuracy on a specific task (a concept known as intrinsic dimension). Therefore, the change in the parameters of a model when fine-tuning may also have a low intrinsic dimension. A reparameterization such as LoRA may achieve competitive results in LLMs.
Equation LABEL:eq:lora-base describes how the LoRA technique is added into a DL module. Given a pre-trained model with $W^{model}\in\mathcal{R}^{d_{m}\times k}$ weights, the change needed for fine-tuning it to a certain task is given by $\Delta W$. Based on the previous hypothesis, it must have a low intrinsic dimension. This is ensured by expressing $\Delta W$ as a low-rank decomposition of two matrices $A$ and $B$, where $A\in\mathcal{R}^{r\times k}$ and $B\in\mathcal{R}^{d_{model}\times r}$. $r$ is the hyperparameter responsible for defining the maximum desired rank of these matrices, therefore it must be $r\ll min(k,d)$. It should be noted that $W^{model}$ is not changing during the fine-tuning phase, and only $A$ and $B$ are trainable. The matrix $B$ is initialized to 0 to ensure that $\Delta W=0$ at the beginning of the training.
@@
LoRA can be applied to almost every module of LLaMA-1 7B, but it is only applied over the queries and values projections, as illustrated in figure 2.
@@
Since LLIAM is based on the LLaMA-1 7B model, most of its hyperparameters are identical to those detailed in [6]. Table 2 lists the main hyperparameters of LLIAM. #heads, #embd, and #decoders refer to the base model architecture, indicating the number of heads in the MHA, the embedding dimension, and the number of stacked decoders, respectively. The remaining hyperparameters relate to the LoRA architecture and training. With $r$ set to 8 and $\alpha$ set to 16, the scale factor when merging the model and LoRA outputs is 2. A light dropout (0.05) is applied to the LoRA weights to prevent overfitting. The optimizer algorithm used is AdamW [51]. A micro-batch training strategy is adopted and a maximum number of training iterations is also defined. To train the LoRAs a supervised procedure is followed. The generated prompt is tokenized and transformed into a nominal representation using the SentencePieces method [52]. Then, input and target tokens are separated. The target tokens logits and the model generated logits are used to calculate a cross-entropy loss.
@@ 4 Experimentation framework
In this section the design of the experiments done to evaluate our proposal is introduced. The datasets, pre-processing pipeline and metrics employed on the experimentation are described in Subsection 4.1. For clarity, the analysis has been divided in two stages, each designed to demonstrate the capabilities of our LLM-based proposal. The two groups are described below:
@@
    Subsection 4.2 presents a comparative study between our proposal, LLIAM, and a set of traditional DL techniques used in TSF, such as RNNs and TCNs, and TimeLLM, a state-of-the-art LLM-based approach. Non-parametric statistical tests, such as Friedman and Wilcoxon, are also used to determine the existence of significant differences between all these algorithms.
@@
    Subsection 4.3 shows a zero-shot study between our proposal and a LLM model. It aims to illustrate how the incorporation of LoRAs and the usage of the described prompting scheme enhances the performance of LLMs as time-series forecasters. The datasets used in this experiment are unknown to both models. This is important as we are evaluating the generalization capacity of the two methods, exploiting their ability to face unknown domains and distributions of the data enabled by Transfer Learning.
4.1 Experimental setup
For the comparative study the Monash University time series forecasting repository has been chosen for selecting almost every dataset used in this study [53]. Of all the available sets, we filtered out all those where the prediction horizon was not clearly specified in the repository and were not uni-variate. Subsequently, we were left with only 7 main sets from this repository (Electricity, M1 Monthly, M1 Quarterly, M3 Monthly, M3 Quarterly, NN5 Daily, NN5 Weekly). In addition, the Weather and ILI datasets has been included, extracted from [37]. They have been widely used as benchmark for the time-series forecasting task. We provide the description of each of the 9 datasets selected below:
  • 1.
    Electricity (1 Dataset): Contains the electricity consumption of 370 customers in kilowatts from 2011 to 2014. All series have the same length. The version whose sampling frequency is weekly has been selected.
  • 2.
    M1 (2 Datasets): This dataset is composed of a selection of series from one of the first competitions dedicated to its prediction, called M competitions (M, because its creator is called Makridakis). It consists of 617 monthly time series, related to economics, industry, and demography. Not all series have the same number of measurements. From it, he versions with monthly and quarterly sampling frequency have been selected.
@@ 4.1 Experimental setup
    M3 (2 Datasets): Like the M1 dataset, it contains 1428 series from the third edition of the M competitions. They are again not of the same length and add finance-related phenomena to their domains. From it, the versions that present a monthly and quarterly sampling frequency have been selected.
@@
    NN5 (2 Datasets): Dataset from another competition, with 111 series collecting transactions from ATMs located in Great Britain. In this case, the series have the same length. From it, the versions with a daily and weekly sampling frequency have been selected.
@@
    Weather (1 Dataset): The Weather dataset [37] includes data collected over a year from 21 meteorological stations in Germany, recorded at 10-minute intervals.
@@
    ILI (1 Dataset): This dataset contains weekly data on individuals exhibiting symptoms similar to Influenza from 2002 to 2020. The data has been gathered via the U.S. Outpatient Influenza-like Illness Surveillance Network (ILINet) [37]. In this dataset, the primary variable of interest is the total number of patients.
For the zero-shot experiment two of the Electricity Transformer Temperature (ETT) datasets [54] and the San Francisco Traffic dataset, both extracted from the Monash University time series forecasting repository are considered. The description of these datasets is presented below:
  • 1.
    San Francisco Traffic (1 Dataset): It contains 862 series collecting occupancy rates of freeways located in the San Francisco Bay Area between the years 2015 and 2016. The version whose sampling frequency is weekly has been selected.
  • 2.
@@
electricity 321 weekly 8 65 Comparative study
m1 617 monthly 18 15 Comparative study
m1 203 quarterly 8 5 Comparative study
m3 1428 monthly 18 15 Comparative study
m3 756 quarterly 8 5 Comparative study
nn5 111 daily 56 9 Comparative study
nn5 111 weekly 8 65 Comparative study
weather 1 10 minutes 96 512 - Comparative study
ILI 1 weekly 24 96 - Comparative study
@@
A sliding window approach is used to adapt the TSF problem as a supervised learning problem. An instance consists of two windows of observations, ordered in time and without gaps. The first represents the input values to be taken by the model ($X$) and the second the expected output ($Y$). Thanks to this restructuring, series has been transformed into a set of $(x,y)$ pairs compatible with LLIAM. With this conversion, a textual prompt is constructed following the template described in Section 3, leaving the datasets ready to be used by the model. Table 3 shows the input window size and the prediction horizon used to construct the sliding windows for each data set, other relevant properties, and in which study each dataset is used.
@@
RNN-LSTM 50, 100, 200 1, 2, 4 32, 64, 128 16, 32 0.001, 0.01 0
RNN-GRU 50, 100, 200 1, 2, 4 32, 64, 128 16, 32 0.001, 0.01 0
TCN 50, 100, 200 1, 2, 4 16, 32, 64 16, 32 0.001, 0.01 0, 0.1, 0.25
Table 4: Hyperparameters to be used in the experimentation for each non-foundational model. dim_h refers to the size of the hidden state in the recurrent networks and the number of channels generated by the convolutions in the TCN. bs indicates the batch size and lr the learning rate of the optimizer.
@@
The evaluation of the models is performed offline. The metrics selected are the Root Mean Square Error [55] (RMSE, equation 5) and the Symmetric Mean Absolute Percentage Error [55] (SMAPE, equation 6). This metric is more suitable for this type of problem because it allows us an easier interpretability of the results by taking into account the magnitude of the data and providing a relative measure unaffected by it. This allows its aggregation and use in statistical tests.
@@
$SMAPE=\frac{2}{n}\sum_{t=0}^{n}\frac{|y_{t}-\hat{y_{t}}|}{|y_{t}|+|\hat{y_{t}}|}$ (6)